\begin{thebibliography}{DFG{\etalchar{+}}11}

\bibitem[Bar96]{barth1996two}
Wolf~Paul Barth.
\newblock Two projective surfaces with many nodes, admitting the symmetries of
  the icosahedron.
\newblock {\em Journal of Algebraic Geometry}, 5(1):173--186, 1996.

\bibitem[Bas06]{basset1906maximum}
Alfred~Barnard Basset.
\newblock The maximum number of double points on a surface.
\newblock {\em Nature}, 73(1889):246, 1906.

\bibitem[BB24]{bouyukliev2024dual}
Iliya Bouyukliev and Stefka Bouyuklieva.
\newblock Dual transform and projective self-dual codes.
\newblock {\em Advances in Mathematics of Communications}, 18:328--341, 2024.

\bibitem[BBK21]{bouyukliev2020computer}
Iliya Bouyukliev, Stefka Bouyuklieva, and Sascha Kurz.
\newblock Computer classification of linear codes.
\newblock {\em IEEE Transactions on Information Theory}, 67:7807--7814, 2021.

\bibitem[Bea79]{beauville1979nombre}
Arnaud Beauville.
\newblock Sur le nombre maximum de points doubles d’une surface dans ${P}^3$
  ($\mu (5)= 31$).
\newblock {\em Journ\'ees de G\'eom\'etrie alg\'ebraique d'Angers}, pages
  207--215, 1979.

\bibitem[BJV00]{bouyukhev2000smallest}
Iliya Bouyukliev, David~Benjamin Jaffe, and Vesselin Vavrek.
\newblock The smallest length of eight-dimensional binary linear codes with
  prescribed minimum distance.
\newblock {\em IEEE Transactions on Information Theory}, 46(4):1539--1544,
  2000.

\bibitem[BR07]{barth2007cusps}
Wolf~Paul Barth and S{\l}awomir Rams.
\newblock Cusps and codes.
\newblock {\em Mathematische Nachrichten}, 280(1-2):50--59, 2007.

\bibitem[Cat81]{catanese1981babbage}
Fabrizio Catanese.
\newblock Babbage's conjecture, contact of surfaces, symmetric determinantal
  varieties and applications.
\newblock {\em Inventiones Mathematicae}, 63(3):433--465, 1981.

\bibitem[Cat96]{catanese1996generalized}
Fabrizio Catanese.
\newblock Generalized {K}ummer surfaces and differentiable invariants of
  {N}oether-{H}orikawa surfaces {I}.
\newblock In Eugenio Calabi, editor, {\em Manifolds and Geometry}, volume~36 of
  {\em Symposia Mathematica}, pages 132--177. Cambridge University Press,
  Cambridge, 1996.

\bibitem[Cay69]{cayley1869vii}
Arthur Cayley.
\newblock A memoir on cubic surfaces.
\newblock {\em Philosophical Transactions of the Royal Society of London},
  159:231--326, 1869.

\bibitem[CC82]{catanese1982constructing}
Fabrizio Catanese and Giuseppe Ceresa.
\newblock Constructing sextic surfaces with a given number d of nodes.
\newblock {\em Journal of Pure and Applied Algebra}, 23(1):1--12, 1982.

\bibitem[CC97]{casnati1997even}
Gianfranco Casnati and Fabrizio Catanese.
\newblock Even sets of nodes are bundle symmetric.
\newblock {\em Journal of Differential Geometry}, 47(2):237--256, 1997.

\bibitem[CC98]{casnati1998errors}
Gianfranco Casnati and Fabrizio Catanese.
\newblock Errors in the paper:``{E}ven sets of nodes are bundle symmetric''.
\newblock {\em Journal of Differential Geometry}, 50(2):415--415, 1998.

\bibitem[CCF{\etalchar{+}}22]{catanese2022varieties}
Fabrizio Catanese, Stephen Coughlan, Davide Frapporti, Alessandro Verra, and
  Michael Kiermaier.
\newblock Varieties of nodal surfaces, coding theory and discriminants of cubic
  hypersurfaces.
\newblock {\em arXiv preprint 2206.05492}, 2022.

\bibitem[CT07]{catanese2007even}
Fabrizio Catanese and Fabio Tonoli.
\newblock Even sets of nodes on sextic surfaces.
\newblock {\em Journal of the European Mathematical Society}, 9(4):705--737,
  2007.

\bibitem[DFG{\etalchar{+}}11]{doran2011codes}
Charles~F. Doran, Michael~G. Faux, Sylvester~James Gates, Tristan H{\"u}bsch,
  Kevin~M. Iga, Gregory~D. Landweber, and Robert~L. Miller.
\newblock Codes and supersymmetry in one dimension.
\newblock {\em Advances in Theoretical and Mathematical Physics},
  15(6):1909--1970, 2011.

\bibitem[DG75]{delsarte1975alternating}
Philippe Delsarte and Jean-Marie Goethals.
\newblock Alternating bilinear forms over $\operatorname{GF}(q)$.
\newblock {\em Journal of Combinatorial Theory, Series A}, 19(1):26--50, 1975.

\bibitem[DGS99]{dodunekov1999some}
Stefan Dodunekov, Sugi Guritman, and Juriaan Simonis.
\newblock Some new results on the minimum length of binary linear codes of
  dimension nine.
\newblock {\em IEEE Transactions on Information Theory}, 45(7):2543--2546,
  1999.

\bibitem[DS98]{dodunekov1998codes}
Stefan Dodunekov and Juriaan Simonis.
\newblock Codes and projective multisets.
\newblock {\em The Electronic Journal of Combinatorics}, 5:1--23, 1998.

\bibitem[End98]{endrass1997minimal}
Stephan Endra\ss.
\newblock Minimal even sets of nodes.
\newblock {\em Journal f\"ur die reine und angewandte Mathematik (Crelles
  Journal)}, 503:87--108, 1998.
\newblock arXiv preprint alg-geom/9710025.

\bibitem[Gab96]{gaborit1996mass}
Philippe Gaborit.
\newblock Mass formulas for self-dual codes over $\mathbb{Z}_4$ and
  $\mathbb{F}_q+u\mathbb{F}_q$.
\newblock {\em IEEE Transactions on Information Theory}, 42(4):1222--1228,
  1996.

\bibitem[HKK18]{honold2018partial}
Thomas Honold, Michael Kiermaier, and Sascha Kurz.
\newblock Partial spreads and vector space partitions.
\newblock In Marcus Greferath, Mario~Osvin Pav{\v{c}}evi{\'c}, Natalia
  Silberstein, and Mar{\'\i}a~{\'A}ngeles V{\'a}zquez-Castro, editors, {\em
  Network Coding and Subspace Designs}, pages 131--170. Springer, 2018.

\bibitem[Jaf00]{jaffe2000optimal}
David~Benjamin Jaffe.
\newblock Optimal binary linear codes of length $\le 30$.
\newblock {\em Discrete Mathematics}, 223(1-3):135--155, 2000.

\bibitem[JR97]{jaffe1997sextic}
David~Benjamin Jaffe and Daniel Ruberman.
\newblock A sextic surface cannot have $66$ nodes.
\newblock {\em Journal of Algebraic Geometry}, 6(1):151--168, 1997.

\bibitem[JT18]{jungnickel2018classification}
Dieter Jungnickel and Vladimir~D. Tonchev.
\newblock The classification of antipodal two-weight linear codes.
\newblock {\em Finite Fields and Their Applications}, 50:372--381, 2018.

\bibitem[KK20]{kiermaier2020lengths}
Michael Kiermaier and Sascha Kurz.
\newblock On the lengths of divisible codes.
\newblock {\em IEEE Transactions on Information Theory}, 66(7):4051--4060,
  2020.

\bibitem[KK23]{kiermaier2020classification}
Michael Kiermaier and Sascha Kurz.
\newblock Classification of $\delta$-divisible linear codes spanned by
  codewords of weight $\delta$.
\newblock {\em IEEE Transactions on Information Theory}, 69(6):3544--3551,
  2023.

\bibitem[Kum64]{kummer1864flachen}
Ernst~Eduard Kummer.
\newblock {\"U}ber die {F}l{\"a}chen vierten {G}rades mit sechzehn
  singul{\"a}ren {P}unkten.
\newblock {\em Collected papers II}, pages 418--432, 1864.

\bibitem[Kur20]{kurz2020classification}
Sascha Kurz.
\newblock Classification of $8$-divisible binary linear codes with minimum
  distance $24$.
\newblock {\em arXiv preprint 2012.06163}, 2020.

\bibitem[Kur24]{kurz2024computer}
Sascha Kurz.
\newblock Computer classification of linear codes based on lattice point
  enumeration.
\newblock In {\em International Congress on Mathematical Software}, pages
  97--105. Springer, 2024.

\bibitem[Lab05]{PhdLabs}
Oliver Labs.
\newblock {\em Hypersurfaces with many singularities -- History, Constructions,
  Algorithms, Visualization}.
\newblock PhD thesis, Mainz, Germany, 2005.

\bibitem[Lab06]{labs2004septic}
Oliver Labs.
\newblock A septic with $99$ real nodes.
\newblock {\em Rendiconti del Seminario Matematico della Università di
  Padova}, 116:299--313, 2006.
\newblock arXiv preprint math/0409348.

\bibitem[Mac62]{macwilliams1962combinatorial}
Florence~Jessie MacWilliams.
\newblock {\em Combinatorial problems of elementary abelian groups}.
\newblock PhD thesis, Radcliffe College, 1962.

\bibitem[Mac63]{macwilliams1963theorem}
Florence~Jessie MacWilliams.
\newblock A theorem on the distribution of weights in a systematic code.
\newblock {\em Bell System Technical Journal}, 42(1):79--94, 1963.

\bibitem[Mil]{MillerWP}
Robert~L. Miller.
\newblock Doubly-even codes.
\newblock \url{https://rlmill.github.io/de_codes/}.
\newblock Accessed: 2025-04-11.

\bibitem[Pet98]{PhdPettersen}
Kjell~Fredrik Pettersen.
\newblock {\em On nodal determinantal quartic hyperfurfaces in $\mathbb{P}^4$}.
\newblock PhD thesis, Oslo, Norway, 1998.

\bibitem[Ple63]{pless1963power}
Vera Pless.
\newblock Power moment identities on weight distributions in error correcting
  codes.
\newblock {\em Information and Control}, 6(2):147--152, 1963.

\bibitem[PT09]{pignatelli2007wahl}
Roberto Pignatelli and Fabio Tonoli.
\newblock On {W}ahl's proof of $\mu(6)= 65$.
\newblock {\em Asian Journal of Mathematics}, 13(3):307--310, 2009.

\bibitem[Sch63]{schlafli1863x}
Ludwig Schl{\"a}fli.
\newblock On the distribution of surfaces of the third order into species, in
  reference to the absence or presence of singular points, and the reality of
  their lines.
\newblock {\em Philosophical Transactions of the Royal Society of London},
  (153):193--241, 1863.

\bibitem[Sim95]{simonis1995macwilliams}
Juriaan Simonis.
\newblock Mac{W}illiams identities and coordinate partitions.
\newblock {\em Linear Algebra and its Applications}, 216:81--91, 1995.

\bibitem[Var83]{varchenko1983semicontinuity}
Alexander~Nicholaevich Varchenko.
\newblock Semicontinuity of the spectrum and an upper bound for the number of
  singular points of the projective hypersurface.
\newblock In {\em Doklady Akademii Nauk SSSR}, volume 270, pages 1294--1297,
  1983.

\bibitem[VGZ18]{van2018genus}
Bert Van~Geemen and Yan Zhao.
\newblock Genus three curves and $56$ nodal sextic surfaces.
\newblock {\em Journal of Algebraic Geometry}, 27(4):583--592, 2018.

\bibitem[Yan97]{yang1997enumeration}
Jin-Gen Yang.
\newblock Enumeration of combinations of rational double points on quartic
  surfaces.
\newblock {\em AMS/IP Studies in Advanced Mathematis}, 5:275--312, 1997.

\end{thebibliography}

\documentclass[notitlepage,12pt]{article}

\usepackage{amsmath}
\usepackage{amsthm}
\usepackage{amssymb}
\usepackage[margin=2cm]{geometry}
\usepackage{url}

\theoremstyle{plain}
\newtheorem{Theorem}{Theorem}

\newtheorem{Corollary}[Theorem]{Corollary}
\newtheorem{Lemma}[Theorem]{Lemma}

\title{Nodal surfaces in $\mathbb{P}^3$ and coding theory}

\author{Sascha Kurz\\ {\small Dept.\ of Mathematics, University of Bayreuth, 95440 Bayreuth, Germany}\\ {\small sascha.kurz@uni-bayreuth.de}}

\begin{document}

\maketitle

\begin{abstract}
\noindent
To each nodal hypersurface one can associate a binary linear code. Here we show that the binary linear 
code associated to sextics in $\mathbb{P}^3$ with the maximum number of $65$ nodes, as e.g.\ the 
Barth sextic, is unique. We also state possible candidates for codes that might be associated with a 
hypothetical septic attaining the currently best known upper bound for the maximum number of nodes.

\smallskip

\noindent
\textbf{Keywords:} Nodal hypersurface, linear code, Barth sextic, coding theory.

\noindent
\textbf{Mathematics Subject Classification:} 14J70, 94B05
\end{abstract}


\section{Introduction}
\label{sec_introduction}
An irreducible hypersurface $S$ of degree $s$ in a complex projective space 
$\mathbb{P}^n:=\mathbb{P}^n(\mathbb{C})$ is the zero set of an irreducible homogeneous 
polynomial $f(x_0,\dots,x_n)$. A singularity is a point on the hypersurface 
where all partial derivatives vanish. A generic hypersurface is smooth i.e.\ it does not
contain any singularities, so that it is natural to ask which combinations of singularities 
can occur on a hypersurface of given degree $s$ in $\mathbb{P}^n$. For $s=1$ there cannot 
be any singularity and for $s=2$ there can be at most one isolated singularity which has 
to be an ordinary double point, or node for brevity, i.e.\ a singularity where the 
Hessian matrix is invertible. The possible combinations of singularities of cubics in $\mathbb{P}^3$
have been classified by Schl\"afli in 1863 \cite{schlafli1863x} and  the classification of all 
quartic surfaces in $\mathbb{P}^3$ with respect to their occurring singularities has been completed
in 1997 \cite{yang1997enumeration}. In $\mathbb{P}^2$ the maximum number of isolated singularities 
of a plane curve of degree $s$ is given by ${s\choose 2}$ and attained by $s$ lines in general position. In 
higher dimensions no such result is known yet. As the classification problem seems to be quite complex for 
larger parameters it makes sense to add further restrictions. A hypersurface is called 
nodal if all of its singularities are nodes. So, let $\mu(s)$ denote the maximum number of nodes of a 
nodal surface in $\mathbb{P}^3$ that can be described by a polynomial of degree $s$. Clearly we have 
$\mu(1)=0$ and $\mu(2)=1$. The values $\mu(3)=4$ and $\mu(4)=16$ are attained by the Cayley cubic \cite{cayley1869vii} and 
a Kummer surface \cite{kummer1864flachen}, respectively. In 1979 Beauville applied coding theory to determine $\mu(5)=31$ 
\cite{beauville1979nombre} and thereby improve a general upper bound of Basset \cite{basset1906maximum}.   
To this end let $\pi\colon \tilde{S} \rightarrow S$ be a minimal resolution 
of singularities. A set $N$ of nodes on $S$ is called even if there exists a divisor $Q$ on $\tilde{S}$ 
such that $2 Q\sim \pi^{-1}(N)$. The even sets of nodes on $S$ comprise the codewords of a binary linear 
code $C$, which we call the code associated to $S$. In \cite{jaffe1997sextic} Jaffe and Ruberman showed 
$\mu(6)<66$ by excluding the existence of certain binary linear codes, so that $\mu(6)=65$ due to the existence of the 
Barth sextic \cite{barth1996two}. As mentioned by Jaffe and Ruberman it would be very interesting to exactly describe 
the associated code for the known examples of nodal surfaces with many nodes. For quintic surfaces this has been 
accomplished by Beauville \cite{beauville1979nombre} and for generalized Kummer surfaces by Catanese \cite{catanese1996generalized}. 
While an explicit equation of the Barth sextic was known its associated code was first determined by showing the 
uniqueness of the possible codes of sextics with $65$ nodes \cite{kurz2020classification}, cf.\ \cite{catanese2022varieties} 
for a more direct derivation. Due to the importance of the problem, since currently none of the two mentioned preprints 
is published, and in order to popularize the problem we would like to give a streamlined proof of the uniqueness of the 
code associated to a nodal sextic with $65$ nodes in this paper. Additionally we give some information on the next open case of 
degree $d=7$.

The remaining part of this paper is structured as follows. In Section~\ref{sec_introduction} we introduce the necessary preliminaries 
from algebraic geometry and coding theory. The uniqueness of the code associated to a sextic in $\mathbb{P}^3$ with the maximum number 
of nodes is then concluded in Section~\ref{sec_uniqueness}. We close with a conclusion and a few remarks on further open problems, 
including the maximum number of nodes of septics in $\mathbb{P}^3$, in Section~\ref{sec_conclusion}.

\section{Preliminaries}
\label{sec_preliminaries}
An $[n,k]_2$ code $C$, or binary linear code with length $n$ and dimension $k$, is a $k$-dimensional subspace of $\mathbb{F}_2^n$. 
The elements of $C$ are called codewords. The dual code $C^\perp$ of $C$ is the subspace that is perpendicular to $C$, i.e.\ $C^\perp$ is 
an $[n,n-k]_2$ code. The weight $\operatorname{wt}(c)$ of a codeword $c\in C$ is the number of non-zero entries in 
$c$ and the minimum weight $d$ of $C$ is the minimum weight over of the non-zero codewords in $C$, so that we also speak of an $[n,k,d]_2$ code. If 
the minimum weight $d^\perp$ of the dual code $C^\perp$ is at least $3$ we say that $C$ is projective. If 
the non-zero weights of codewords in $C$ are contained in $\left\{w_1,\dots,w_l\right\}$, then we say that $C$ is a $\Delta$-divisible 
$[n,k,\{w_1,\dots,w_l\}]_2$ code if $w_i$ is divisible by $\Delta$ for all $1\le i\le l$. The polynomial $W_C(x):=\sum_{c\in C} x^{\operatorname{wt}(c)}y^{n-\operatorname{wt}(c)}$ is 
called the weight enumerator of $C$. The MacWilliams identity determines the weigh enumerator of the dual code via $W_{C^\perp}(x,y)=\tfrac{1}{2^k}\cdot W_C(x+y,x-y)$ 
\cite{macwilliams1962combinatorial,macwilliams1963theorem}. 
Each $k\times n$ matrix over $\mathbb{F}_2$ whose row span equals $C$ is called a generator matrix of $C$. If 
$C$ contains a codeword with a non-zero entry in position $i$ for all $1\le i\le n$, which is equivalent to $d^\perp\ge 2$, then $C$ has full length and $n$ is called the effective length 
of $C$. In order to ease the notation we only consider $[n,k]_2$ codes of full length. For a given codeword $c$ of an $[n,k]_2$ code the support 
$\operatorname{supp}(c)$ is given by $\{1\le i\le n\,:\, c_i\neq 0\}$, i.e., the cardinality of its support equals its weight. With this the residual 
code $\operatorname{Res}(C;c)$ of an $[n,k]_2$ code $C$ with respect to a non-zero codeword $c$ is the restriction of the codewords to 
$\{1,\dots,n\}\backslash \operatorname{supp}(c)$, which has effective length $n-\operatorname{wt}(c)$.

The code associated to the Cayley cubic is an $[4,1,\{4\}]_2$ code with weight enumerator $W_C(x,y)=x^0y^4+x^4y^0$ and the code associated with a Kummer surface 
is an $[16,5,\{8,16\}]_2$ code with weight enumerator $W_C(x,y)=x^0y^{16}+30x^8y^8+x^{16}y^0$. Geometrically the latter code corresponds to the $16$ points of an affine 
solid and a generator matrix is given by
$$
  \left(\begin{smallmatrix}
  0000000011111111\\
  0000111100001111\\
  0011001100110011\\
  0101010101010101\\
  1111111111111111
  \end{smallmatrix}\right).
$$ 
The code associated to a quintic with $31$ nodes is given by a $[31,5,\{16\}]_2$ code with weight enumerator $W_C(x,y)=x^0y^{31}+31x^{16}y^{15}$. Geometrically such a 
code corresponds to the $31$ points of the projective geometry $\operatorname{PG}(4,2)$ and a generator matrix is given by
$$
  \left(\begin{smallmatrix}
  0000000111111110000000011111111\\
  0001111000011110000111100001111\\
  0110011001100110011001100110011\\
  1010101010101010101010101010101\\
  0000000000000001111111111111111
  \end{smallmatrix}\right).
$$
The following general properties of the associated $[n,k]_2$ code $C$ of a nodal surface with degree $s$ and $m$ nodes are known. For the length we have $n\le m$ 
and for the dimension a general argument of Beauville \cite{beauville1979nombre} gives $k\ge m-\left\lceil s^3/2\right\rceil +2s^2-3s+1$, see 
\cite[Proposition 4.3]{jaffe1997sextic}. If $d$ is odd, then $C$ is $4$-divisible and $8$-divisible otherwise, see \cite[Proposition 2.11]{catanese1981babbage}. 
The minimum distance $d$ satisfies $d\ge 2\lceil s(s-2)/2\rceil$, see \cite[Theorem 1.10]{endrass1997minimal}. So, for sextics with $65$ nodes the associated code 
is a $8$-divisible $[n,k,24]_2$ code with $n\le 65$ and $k\ge 12$.

\section{Uniqueness of the code associated to a sextic with $\mathbf{65}$ nodes}
\label{sec_uniqueness}

First we want to conclude some further properties of $8$-divisible $[n,12,24]_2$ codes (of full length) by purely coding theoretic arguments. First we observe that 
for a $q^r$-divisible linear code over $\mathbb{F}_q$ with $r\ge 1$ each residual code is $q^{r-1}$-divisible, see e.g.\ \cite[Lemma 7]{honold2018partial}. Note 
that we have $C\subseteq C^\perp$ for each $4$-divisible $[n,k]_2$ code, so that $k\le n/2$. Setting $W_C(x,y)=\sum_i a_ix^iy^{n-i}$ and $W_{C^\perp}(x,y)=\sum_i a_i^*x^iy^{n-i}$ 
the equations for the coefficients $y^0$, $y^1$, $y^2$, and $y^3$ in the MacWilliams identity can be rewritten to 
\begin{eqnarray}
  \sum_{i>0} a_i &=& 2^k-1,\\
  \sum_{i\ge 0} ia_i &=& 2^{k-1}n,\\
  \sum_{i\ge 0} i^2a_i &=& 2^{k-1}(a_2^*+n(n+1)/2),\\ 
  \sum_{i\ge 0} i^3a_i &=& 2^{k-2}(3(a_2^*n-a_3^*)+n^2(n+3)/2).
\end{eqnarray} 
We also speak of the first four MacWilliams identities. In this special form, those equations are also known as the 
first four (Pless) power moments \cite{pless1963power}.

\begin{Lemma}
  \label{lemma_n_ge_63}
  Let $C$ be a binary $8$-divisible linear code with minimum distance $d\ge 24$, dimension $k=12$ and effective length $n\le 65$, then
  $a_{40}\ge 1$ and $n\ge 63$.
\end{Lemma}
\begin{proof}
  Solving the first four MacWilliams identities for $a_{24}$, $a_{32}$, $a_{40}$, and $a_{48}$ gives
  $$
    a_{40}=\frac{205}{2}n^2-6808n-\frac{1}{2}n^3+(208-3n)a_2^*+3a_3^*+6a_{56}+20a_{64}+147420 
  $$
  and
  $$
    a_{40}+a_{48}=71n^2-\frac{14504}{3}n-\frac{1}{3}n^3+(144-2n)a_2^*+2a_3^*+2a_{56}+10a_{64}+106470.
  $$    
  Since $a_2^*,a_3^*,a_{56},a_{64}\ge 0$, $208-3n\ge 0$, $144-2n\ge 0$ we have 
  $$
    a_{40}\ge \frac{205}{2}n^2-6808n-\frac{1}{2}n^3+147420  
  $$
  and   
  $$
    a_{40}+a_{48}\ge 71n^2-\frac{14504}{3}n-\frac{1}{3}n^3+106470.  
  $$  
  For $54\le n\le 60$ we have $a_{40}+a_{48}<0$, which is impossible. If either $n\le 53$ or $61\le n\le 65$, 
  then $a_{40}\ge 1$. Thus, $a_{40}\ge 1$. Consider the residual code $C'$ of a codeword of weight $40$. $C'$ has dimension $11$ 
  and is $4$-divisible, so that its length is at least $23$, see e.g.\ \cite[Section VIII]{gaborit1996mass}.  
\end{proof}

We remark that there even exists an $8$-divisible $[64,13,24]_2$ code $C$
with generator matrix 
$$
\left(\begin{smallmatrix}
1111111111111111111111111111111111111111111111111111111111111111\\
0000000000000000000000000000000011111111111111111111111111111111\\
0000000000000000111111111111111100000000000000001111111111111111\\
0000000011111111000000001111111100000000111111110000000011111111\\
0000111100001111000011110000111100001111000011110000111100001111\\
0011001100110011001100110011001100110011001100110011001100110011\\
0101010101010101010101010101010101010101010101010101010101010101\\
0000000000001111001100110011110000111100001100111111000011111111\\
0001001000011101000100100001110101001000010001110100100001000111\\
0000011001100000001110101010001101011100110001011001111111111001\\
0000000001101001000000001001011001011010110011000101101000110011\\
0001001000011101011110110111010000101110110111101011100001001000\\
0000001101010110011001011100111100000011101010011001101011001111\\
\end{smallmatrix}\right).
$$
With these properties the code is unique, as shown by exhaustive enumeration in \cite{kurz2020classification}.  
Its automorphism group has order $23224320$ and its weight enumerator 
is given by $W_C(x,y)=x^{0}y^{64}+1008x^{24}y^{40}+6174x^{32}y^{32}+1008x^{40}y^{24}+x^{64}y^{0}$. 
The code was obtained in \cite{delsarte1975alternating} and has the following nice description, see \cite{jaffe1997sextic}:
It is a subcode of the second order Reed-Muller code $R(2,6)$ containing the first order Reed-Muller code $R(1,6)$ as a subcode. The
cosets of $R(1,6)$ in it correspond to the symplectic forms $B_a$ in $\mathbb{F}_{64}$, given by $B_a(x,y)=\operatorname{tr}((ax^4 + a^{16}x^{16})y)$.
So, subcodes of this code give examples of $8$-divisible $[n,12,24]_2$ codes for $n=63$ and $n=64$. We remark that an $[59,12,24]_2$ code exists 
while the existence of a $[58,12,24]_2$ code is currently unknown. So the $8$-divisibility of the code is a severe restriction.

\begin{Lemma}
  Let $C$ be an $8$-divisible $[65,12,24]_2$ code. Then $C$ does not contain a codeword of weight $64$ or $56$.
\end{Lemma}
\begin{proof}
  The residual code of a codeword of weight $64$ would be a $4$-divisible code of length $1$, which obviously cannot exist. Similarly, the residual 
  code of a codeword of weight $56$ would be a $4$-divisible code of length $9$. However, no such code exists \cite{kiermaier2020lengths}.
\end{proof}

We remark that it is also possible to exclude the existence of a codeword of weight $48$ in an $8$-divisible $[65,12,24]_2$ code $C$ by theoretical arguments. 
For a codeword $c$ of weight $48$ the residual code $\operatorname{Res}(C;c)$ is a $[17,k,\{4,8,12\}]_2$ code, so that $k\le 8$ \cite[Section VIII]{gaborit1996mass}. The restriction of the 
codewords of $C$ to the coordinates in the support of $c$ gives a $[48,k',\{24,48\}]_2$ code $C'$. Codes whose only occurring non-zero weights are the minimum distance $d$
and $2d$ are completely characterized in \cite{jungnickel2018classification}. In our situation $C'$ is a uniquely defined $[48,k',\{24,48\}]_2$ code with $k'\in\{4,5\}$, 
so that $k\in\{7,8\}$. Up to isomorphism there are only five $[17,7,\{4,8,12\}]_2$ codes and a unique $[17,8,\{4,8,12\}]_2$ code \cite{doran2011codes,MillerWP}. For a 
computer-free classification for the possibilities for $\operatorname{Res}(C;c)$ one may consider the decomposition of $\operatorname{Res}(C;c)$ into subcodes spanned by codewords of weight $4$, which is
completely characterized in a more general setting \cite{kiermaier2020classification}. Having the explicit possible choices for $C'$ and $\operatorname{Res}(C;c)$ at hand one can easily 
exclude the existence of $C$ with a codeword $c$ of weight $48$. However, since those arguments are quite lengthy, due to many case differentiations, and such a result 
can be easily obtained by exhaustive computer enumeration within seconds we do not go into details.    

\begin{Lemma}
  \label{lemma_ten_cases}
  Let $C$ be an $8$-divisible $[n,12,24]_2$ code with $n\le 65$. Then the non-zero weights of $C$ are contained in $\{24,32,40,64\}$.
\end{Lemma}
\begin{proof}
  From Lemma~\ref{lemma_n_ge_63} we know that $n\ge 63$ and that $C$ contains a codeword $c$ of weight $40$. The residual code then is a $4$-divisible $[n-40,11]_2$ code.
  Using the software package \texttt{LinCode} \cite{bouyukliev2020computer} we have exhaustively enumerated all possibilities for $\operatorname{Res}(C;c)$ up to isomorphism.
  There are exactly $11$ $[23,11]_2$, $83$ $[24,11]_2$, and $215$ $[25,11]_2$ $4$-divisible codes cf.\ \cite{doran2011codes,MillerWP}. Starting from these residual codes 
  we have exhaustively enumerated all possibilities for $C$ using \texttt{LinCode}. Up to isomorphism there are unique $[n,12,24]_2$ codes for $n\in\{63,65\}$ and eight 
  $[64,12,24]_2$ codes (assuming $8$-divisibility) and in none of these cases codewords of weight $48$ or $56$ occur. (Generator matrices for these ten cases are given 
  in \cite{kurz2020classification}.) 
\end{proof}

We remark that the initial classification of the $8$-divisible $[\le 65,12,24]_2$ codes in \cite{kurz2020classification} took almost 1000 single-core CPU hours on a computing cluster. 
There are e.g.\ $978528$ $8$-divisible $[63,10,24]_2$ and $704571$ $8$-divisible $[64,11,24]_2$ codes. Constructing the $8$-divisible $[\le 65,12,24]_2$ codes via the residual 
codes of a codeword of weight $40$ allows to keep the numbers of intermediate codes in the enumeration process much smaller. The determination problem of possible 
codes associated with a sextic with $65$ nodes was indeed the initial motivation to develop a new enumeration algorithm for linear codes. Over the time there were algorithmic 
improvements, see \cite{kurz2020classification,kurz2024computer} for details, so that the computation underlying Lemma~\ref{lemma_ten_cases} can now be performed in less than 
two hours on a single core. 

In \cite{jaffe1997sextic} it was shown that the code associated to a sextic with $65$ nodes cannot contain a codeword with weight $48$ or $64$. This boils down the number of possible codes
to three: 
\begin{Lemma}
  \label{lemma_three_cases}
  Let $C$ be an $[n,12,\{24,32,40,48,56\}]_2$ code with $n\le 65$. Then $C$ is isomorphic to one of the three cases:
    \begin{enumerate}
    \item[(1)]
    $[63,12,24]_2$\\
    $\left(\begin{smallmatrix} 
    001100001110000001111101000011111110010010100100001100000000000\\
    101001111111000000110111010000100110100011011000000010000000000\\
    000100111011100011110111001000010000110000110110100001000000000\\
    010001111111110011001100001001100100010001101000001000100000000\\
    110001110000010111001111011000011100100011000010100000010000000\\
    000000011000110111100011010011101110010001011110000000001000000\\
    010011110001111101010000110100100011101110111111111000000100000\\
    001000110111101100001111110000000001100110000111100000000010000\\
    000111110001100011000000001100011111100001100001111000000001000\\
    000000001111100000111111111100000000011111100000011000000000100\\
    000000000000011111111111111100000000000000011111111000000000010\\
    000000000000000000000000000011111111111111111111111000000000001\\
    \end{smallmatrix}\right)$\\
    $W_C(x,y)=x^{0}y^{63}+630x^{24}y^{39}+3087x^{32}y^{31}+378x^{40}y^{23}$\\
    $\# \operatorname{Aut}(C)=362880$
    \item[(2)]
    $[64,12,24]_2$\\
    $\left(\begin{smallmatrix}
    0000110001101110000100100100100011011000011011011110100000000000\\
    1011110000100110010000001100010000111101001110111000010000000000\\
    1010110001001010110010000000101111110000001100101011001000000000\\
    1111100000001100000010100100111101000011011011101000000100000000\\
    0111000000001010110110001100011000000110111100110011000010000000\\
    0000000100001001111110011010010101001101010101010101000001000000\\
    0101011111010000010001111001110011000100100000101100000000100000\\
    0011010011001000001111111001111111011111010011011100000000010000\\
    0000101111000110000000000111101111000011001111000011000000001000\\
    0000011111000001111111111111100000111111000000111111000000000100\\
    0000000000111111111111111111100000000000111111111111000000000010\\
    0000000000000000000000000000011111111111111111111111000000000001\\
    \end{smallmatrix}\right)$\\
    $W_C(x,y)=x^{0}y^{64}+502x^{24}y^{40}+3087c^{32}y^{32}+506x^{40}y^{24}$\\
    $\# \operatorname{Aut}(C)=5760$
    \item[(3)]
    $[65,12,24]_2$\\
    $\left(\begin{smallmatrix}  
    10000100000000110110010001110100111101010001011110010100000000000\\
    10100100011000001001000110100110111111001000001100011010000000000\\
    01000010011100011000000100110100110000011111011110001001000000000\\
    11110100001110110100000011010110100001011100000100001000100000000\\
    01101011000001100011010001000011001010001111000010111000010000000\\
    00101001110111101011000001011000000110111001001000100000001000000\\
    00011000111111100000111110001000100010001010101001100000000100000\\
    00000111001011100111110001010100000001111001100000011000000010000\\
    00011111000111100000001111001101111110000111100111111000000001000\\
    00000000111111100000000000111100011110000000011111111000000000100\\
    00000000000000011111111111111100000001111111111111111000000000010\\
    00000000000000000000000000000011111111111111111111111000000000001\\
    \end{smallmatrix}\right)$\\
    $W_C(x,y)=x^{0}y^{65}+390x^{24}y^{41}+3055x^{32}y^{33}+650x^{40}y^{25}$\\
    $\# \operatorname{Aut}(C)=15600$
  \end{enumerate}
  All three codes are projective.
\end{Lemma}

\begin{Theorem}
  \label{thm_main}
  If $S$ is a sextic with $65$ nodes in $\mathbb{P}^3$ then its associated code is given by Lemma~\ref{lemma_three_cases}.(3).
\end{Theorem}
\begin{proof}
  Let $C$ be the code associated to a sextic in $\mathbb{P}^3$. In \cite{endrass1997minimal} the existence of a code $C'\supsetneq C$ with $\dim(C')=\dim(C)+1$ 
  such that the codewords in $C'\backslash C$ have weights contained in $\{16,28,32,36,\dots\}$ was shown. First we use \texttt{LinCode} to check that none 
  of the cases of Lemma~\ref{lemma_three_cases} can be extended to a $[\le 66,13,\{24,32,40,48,56\}]_2$ code, so that there are just three choices for $C$. Then 
  we used \texttt{LinCode} to extend $C$ to $C'$.   
\end{proof}
We remark that the code $C'$ in the proof of Theorem~\ref{thm_main} is unique. A generator matrix its given by
$$
  \left(\begin{smallmatrix}
  100000000011000100000110100000010001100011111110011111001111001010\\
  010000001000101000001110000010111100000110011011000100101100010100\\
  001000001011111100001011110000101110100110011000100000110110110111\\
  000100001001000100000111100001010001101010101010011000001010001011\\
  000010001010010000001101000101100000111001010100010101000101110010\\
  000001001001111000000101011000111011000100111111010111011001011100\\
  000000101011001100001100100010110101100111111111011110000011001100\\
  000000011010110100000110101100100010101100111110011111001111001010\\
  000000000111100000001111101001010011110011111110010110010101011010\\
  000000000000000010001000001000100100010011111111101110101010100101\\
  000000000000000001000100000100011000100011111111011101101001011010\\
  000000000000000000100010100010000001000111111110111011010110101010\\
  000000000000000000010001010001000010001011111101110111010101010101
  \end{smallmatrix}\right).
$$ 
The automorphism group of $C'$ has order $15600$ and the weight enumerator of $C'\backslash C$ is given by $26x^{16}y^{50} 
+ 650x^{28}y^{38} + 1690x^{32}y^{34} + 1300x^{36}y^{30} + 300x^{40}y^{26} + 130x^{44}y^{22}$. For a nice description of 
the associated code $C$ of a sextic with $65$ nodes and its automorphism group we refer to \cite[Section 4]{catanese2022varieties}. Since the residual codes 
of codewords of weight $40$ played an important role in the determination of $C$ we mention that they all are unique with weight enumerator 
$W_{\operatorname{Res}(C;c)}(x,y)=x^{0}y^{25}+3x^{4}y^{21}+258x^{8}y^{17}+1278x^{12}y^{13}+493x^{16}y^9+15x^{20}y^5$ and an automorphism group of order $4608$. 
A generator matrix is given by 
$$
  \left(\begin{smallmatrix}  
    0000100100100100000000000\\ 
    0100011010100111000000000\\ 
    1010010000101110100000000\\ 
    1100100010101010010000000\\ 
    1111101111001010001000000\\ 
    0101111000010100000100000\\ 
    0011101100010010000010000\\ 
    0011100001101010000001000\\ 
    0000100000000110000000100\\ 
    0000011100011110000000010\\ 
    0000000011111110000000001\\ 
  \end{smallmatrix}\right).
$$

While Theorem~\ref{thm_main} shows the uniqueness of the code associated to a sextic with the maximum number of nodes, the sextic itself is far from being unique. In 
\cite[Theorem 5.5.9]{PhdPettersen} a $3$-parameter family of sextics with $65$ nodes, including the Barth sextic, was obtained; cf.\ \cite[Theorem 196]{catanese2022varieties}.

The non-zero weights occurring in a code associated to a nodal sextic in $\mathbb{P}^3$ are contained in $\{24,32,40,56\}$ and all cases can indeed occur \cite{catanese2007even,van2018genus}.
The exclusion of weight $56$ in \cite{casnati1997even} turned out to be incorrect \cite{casnati1998errors}. In \cite[Lemma 2.3]{pignatelli2007wahl} it was show that a binary code 
with non-zero weights in $\{24,32,56\}$ has dimension at most $10$ and in \cite[Lemma 2.1]{pignatelli2007wahl} it was show that a binary code with non-zero weights in $\{24,32\}$ has dimension 
at most $9$.
\begin{Lemma}
  Let $C$ be a $[n,k,\{24,32\}]_2$ code with $n\le 56$. Then, we have $k\le 8$. Moreover, if $k=8$, then $n\in\{51,54,55,56\}$. For $(n,k)=(51,8)$ the code is projective.
\end{Lemma} 
\begin{proof}
  Since each $[n,8,24]_2$ code satisfies $n\ge 51$, see \cite{bouyukhev2000smallest}, we have $51\le n\le 56$. The fact that each $[n,7,24]_2$ code satisfies $n\ge 50$ implies 
  that the code is projective for $(n,k)=(51,8)$. For an $[n,8,\{24,32\}]_2$ code we apply the dual transform as in \cite[Theorem 3.3]{bouyukliev2024dual} with $\alpha=\tfrac{1}{8}$ and 
  $\beta=-3$. The projective dual code $D_{\alpha,\beta}(C)$ has length $128\alpha n+255\beta$ and minimum distance $64(\alpha n+2\beta)$ if $n<2^k-1$.  
  Since no $[67,8,32]_2$ code and no $[83,8,40]_2$ code exists, see \cite{bouyukhev2000smallest}, we have $n\notin\{52,53\}$. Let $C'$ be the $[56,9,\{24,32,56\}]_2$ code arising from 
  $C$ by adding a codeword of weight $56$. Then we clearly have $a_{56}=1$ and $a_{24}=a_{32}$ for $C'$. Using the MacWilliams identities we compute $a_{24}=a_{32}=255$ and $a_2^*=10$.
  Let $\mathcal{P}'$ be the multiset of points of cardinality $56$ in $\operatorname{PG}(8,2)$ corresponding to $C'$, see e.g.\ \cite{dodunekov1998codes}. Projection of $\mathcal{P}'$ 
  through a point of multiplicity $m\ge 1$ yields a multiset of points $\mathcal{P}$ of cardinality $56-m$ in $\operatorname{PG}(7,2)$ that corresponds to an $[56-m,8,\{24,32\}]_2$ code, 
  so that $m\in\{1,2,5\}$. Since each point of multiplicity $m$ contributes ${m\choose 2}$ to $a_2^*$ the multiset $\mathcal{P}'$ consists either of one point of multiplicity $5$ and $51$ 
  points of multiplicity $1$ or five points of multiplicity $2$ and $46$ points of multiplicity $1$.
  
  Now assume that $H$ is an $[l,9,\{24,32\}]_2$ code and $H'$ the $[56,10,\{24,32,56\}]_2$ code that arises from $H$ by adding a codeword of weight $56$. Let $\mathcal{Q}$ and 
  $\mathcal{Q}'$ be the multisets of points corresponding to $H$ and $H'$, respectively. For $H'$ we compute $a_{24}=a_{32}=511$, $a_{56}=1$, $a_2^*=7$, and $a_3^*=0$. Clearly, 
  $\mathcal{Q}'$ contains a point of multiplicity $1$ and projection through this point yields a multiset of points that corresponds to a $[55,9,\{24,32\}]_2$ code. W.l.o.g.\ we 
  assume that $H$ was chosen as such a code, i.e., we have $l=55$. From the classification of the possible lengths of $[\le 56,8,\{24,32\}]_2$ codes we conclude that the 
  point multiplicities in $\mathcal{Q}$ are contained in $\{0,1,4\}$. From the MacWilliams identities we compute $a_{24}=284$, $a_{32}=227$, and $a_2^*=7$ for $H$. 
  Since each point of multiplicity $m$ contributes ${m\choose 2}$ to $a_2^*$ this is impossible. 
\end{proof}
Up to isomorphism there are unique $[51,8,\{24,32\}]_2$ and $[54,8,\{24,32\}]_2$ codes, two $[55,8,\{24,32\}]_2$ codes, three two $[56,8,\{24,32\}]_2$ codes, and 
two $[56,9,\{24,32,56\}]_2$ codes. Generator matrices are given by
\begin{eqnarray*}
  \!\!\!\!\!\!&&\left(\begin{smallmatrix}
    101100011111111111100001010111111111100111110000000\\
    010011011111111000011110101111111110011111101000000\\
    000000100001111111111110000000011111111001100100000\\
    001111100110011001100110011001100110101010100010000\\
    001111111000011110011001100001111000101100100001000\\
    110111100110000011101011000010111010010101100000100\\
    110101101010101010011100011100001011100110100000010\\
    000110100011001100101101111111010101010100100000001  
  \end{smallmatrix}\right),
  \left(\begin{smallmatrix}
    111111111111000000111111111111110000000001111110000000\\
    000000000000111111111111111111110000000100011001000000\\
    000000011111000000000000001111111111111001111100100000\\
    000111111111111111000011111111110001111010000100010000\\
    001000111111001111001100110000110010001010111000001000\\
    001111000111000001000101110011000110011111100100000100\\
    011111111011000110000110001101001110010010001000000010\\
    101001111001110000000110110001001001101011101100000001
  \end{smallmatrix}\right),\\
  \!\!\!\!\!\!&&\left(\begin{smallmatrix}
    1111111111111000000111111111111110000000001111010000000\\
    0000000000000111111111111111111110000000100011001000000\\
    0000000111111000000000000001111111111111001111000100000\\
    0001111111111111111000011111111110001111010000000010000\\
    0110011001111000011000011110000110110011001100100001000\\
    0011101010001001101011101111111001011111111111000000100\\
    1110001101110011111000110110111010111100111111000000010\\
    1111101010010101100100011010001100001100110011000000001
  \end{smallmatrix}\right),
  \left(\begin{smallmatrix}
    1111111111111111111000000000000000111111111111010000000\\
    1111111110000000000111111111100000111111111110101000000\\
    1111111110000111111000001111111111000000111110100100000\\
    0000000000000000111001111111100111000111111111000010000\\
    0000011110011111011000010001111001111000000011100001000\\
    1111100010100011111010110110100010001011000100000000100\\
    1111100101001101001010000111101000011100001001100000010\\
    0000001110110011010111110010010000001101011111000000001
  \end{smallmatrix}\right),\\
  \!\!\!\!\!\!&&\left(\begin{smallmatrix}
    11111111111111111111111111000000000000111110000010000000\\
    00000011111111111111000000111111000000001101000001000000\\
    11111100000000111111000000000000111111111100100000100000\\
    11111100001111111111001111111111001111000000010000010000\\
    00111100000011000011111111000011110011001100001000001000\\
    11001100111100001100110011001100001111110000000100000100\\
    00011111001100110100000000010101001111010101011100000010\\
    11001101011111001011000100000001010111000111000100000001  
  \end{smallmatrix}\right),
  \left(\begin{smallmatrix}
    11111111111111111111111111111110000000000000000010000000\\
    00000000111111111111111100000001111000011100000001000000\\
    11111111000000001111111100000000000111100011100000100000\\
    00001111000011110000111100000001111111100000011100010000\\
    00111111011100010001000100011110011011100000100100001000\\
    00010001100100110001011100000110100111111101101000000100\\
    11110011101111010111100000011110000101101111111100000010\\
    01011111111000010001011101100001001010001101010000000001
  \end{smallmatrix}\right),\\
  \!\!\!\!\!\!&&\left(\begin{smallmatrix}
    11111111111111111111111111111110000000000000000010000000\\
    00000000111111111111111100000001111000011100000001000000\\
    11111111000000001111111100000000000111100011100000100000\\
    00001111000011110000111100000001111111100000011100010000\\
    00010111001100110001000101111110001001100100011100001000\\
    11111001011100010001011110111110110011101111100100000100\\
    01100000110101000010111010111110011010011000101000000010\\
    10101010000111010111101110111111111000110101011100000001
  \end{smallmatrix}\right),
  \left(\begin{smallmatrix}
    11111111111111111111111111000000000000111110000100000000\\
    00000011111111111111000000111111000000001101000010000000\\
    11111100000000111111000000000000111111111100100001000000\\
    11111100001111111111001111111111001111000000010000100000\\
    00111100000011000011111111000011110011001100001000010000\\
    00000101110100001100011111000111000001011111101000001000\\
    00001100011011010011100000011101111101000110110000000100\\
    11000001110100011111100011001001001110001010110000000010\\
    11110111101000111100101100101111001110101111101000000001  
  \end{smallmatrix}\right),\\
  \!\!\!\!\!\!&&\left(\begin{smallmatrix}
    11111111111111111110000001111100000000001111111100000000\\
    11111111111111000001111110000011111000001111101010000000\\
    00000000000000111111111110000000000111111111111001000000\\
    00000001111111111110001110011111111001111111100000100000\\
    00111110001111011111110110011100111110010011110000010000\\
    01000000010111101110010001100101011110010101110000001000\\
    11000000111011110111100010101100100010110010010000000100\\
    00111110010001000010110101100100001101011110101000000010\\
    01111110110010000101000000101110110110010101001000000001  
  \end{smallmatrix}\right).
\end{eqnarray*}  

\begin{Corollary}
  Let $C$ be a $[56,k,\{24,32,56\}]_2$ code. Then, we have $k\le 9$.
\end{Corollary}

We remark that there exists a $[56,10,24]_2$ code with weight enumerator $W_C(x,y)=x^0y^{56}+399x^{24}y^{32}+224x^{28}y^{28}+399x^{32}y^{24}+x^{56}y^{0}$.

\section{Conclusion and outlook}
\label{sec_conclusion}
We have determined the unique code associated to a sextic in $\mathbb{P}^3$ with the maximum number of nodes in Theorem~\ref{thm_main}. For each $0\le \mu\le 65$ there exists a sextic 
in $\mathbb{P}^3$ with $\mu$ nodes \cite{barth1996two,catanese1982constructing}. Using similar techniques for sextics with $64$ codes there currently remain just seven possible candidates 
for the associated code \cite[Theorem 227]{catanese2022varieties}. However for sextics with $63$ nodes such a list of candidates would be rather huge, so that there is a need for more 
necessary conditions on the associated codes that might by exploited using coding theory techniques. 

For septics in $\mathbb{P}^3$ the currently best known bounds on the maximum number of nodes are given by $99\le \mu(7)\le 104$ \cite{PhdLabs,labs2004septic,varchenko1983semicontinuity}. 
With respect to the upper bound we mention that the $[96,10,44]_2$ code with weight enumerator $W_C(x,y)=x^0y^{96}+504x^{44}y^{52}+124x^{48}y^{48}+336x^{52}y^{44}+56x^{60}y^{36}+3x^{64}y^{32}$ 
given by the generator matrix
$$
  \left(\begin{smallmatrix}
    100000000010000100111101100000110001001101111110001000001101010100110011101101110011111000101100\\
    010000000001100010011110110000011000000110111111100100000110101010011001110110111001011100010110\\
    001000000010010101110010111000111101101110100001011010001110100001111111010110101111110110100111\\
    000100000001001010111001111100011110110111010000001101000111110000111111101011010111111011010011\\
    000010000010100001100001111110111110010110010110101110101110101100101100011000011000000101000101\\
    000001000001110000110000011111011111001011001011010111010111010110010110001100001100100010100010\\
    000000100010011100100101001111011110101000011011100011100110011111111000101011110101101001111101\\
    000000010001101110010010000111101111110100001101010001110011001111111100110101111010110100111110\\
    000000001010010011110100000011000110110111111000100000110100010011001101110111001110100010110011\\
    000000000101001001111010000001100011011011111100010000011010101001100110011011100111110001011001
  \end{smallmatrix}\right)
$$
satisfies all constraints on the weights from \cite{endrass1997minimal}. In the special case where an even set of nodes 
is cut out by a smooth cubic surface the corresponding codeword must have weight $36$, while weight $40$ cannot occur \cite{endrass1997minimal}. To this end,
a $[94,10,36]_2$ code with weight enumerator $W_C(x,y)=x^0y^{94}+120x^{36}y^{58}+182x^{44}y^{50}+489x^{48}y^{46}+192x^{52}y^{42}+14x^{56}y^{38}+26x^{60}y^{34}$ 
is given by the generator matrix
$$
  \left(\begin{smallmatrix}
    1111111111111110000000000001111110000001111111111110000001100000000000000000000000100000000000\\
    0000001111110001111110000000000001111111111111111110001110011000000000000000000000010000000000\\
    0001110001110000000001111111111111111110000001111110001110000110000000000000000000001000000000\\
    1111111111110001111111111110000001111110000000000001110000000001100000000000000000000100000000\\
    0001111110001110001110001110001110001110001110001111111110000000011000000000000000000010000000\\
    0001110000001110000001110001111111110001111110000001111110000000000110011100000000000001000000\\
    0001110001111111111110001111110001110000001110001110000000000000000001111100000000000000100000\\
    0010010010010110110110010110010010110010110010010010010110111110101110101111111000000000010000\\
    1111111110000000001111110000001110000000001111110001110000000000000000011100011111000000001000\\
    0000110100100010011010111010100101010001010110110010110011101011010011100101101011000000000111
  \end{smallmatrix}\right).
$$
So, additional restrictions on the code are needed in order to improve upon the upper bound for $\mu(7)$ using coding theoretic methods. The algorithmic tools for 
the exhaustive enumeration of linear codes are available. A first step might be the computation of associated codes of known hypersurfaces with many nodes.

For another type of singularities, so-called cusps, one can associate a $3$-divisible code over $\mathbb{F}_3$ \cite{barth2007cusps}, see also 
\cite{PhdLabs} for a general overview on hypersurfaces with many singularities.

On the coding theory side there are a few more sophisticated techniques that we have not applied in Section~\ref{sec_uniqueness}. If a binary linear code is 
$2^r$-divisible then there are some restrictions on the number of codewords whose weights are divisible by $2^{r+1}$, see e.g.\ \cite[Proposition 5]{dodunekov1999some}. 
The MacWilliams identities can be generalized to coordinate partitions \cite{simonis1995macwilliams}. In combination with linear programming this was used quite successfully 
to show the non-existence of some binary linear codes or to even show uniqueness of some optimal codes, see e.g.\ \cite{jaffe2000optimal}. However, applying those techniques 
usually comes with extensive computer calculations. So, in order to keep the paper relatively short and since all of the used computer calculations were performed in 
a reasonable short amount of time we have not applied those techniques to our problem.

\newcommand{\etalchar}[1]{$^{#1}$}


\end{document}